# A Communication Consistent Approach to Signal Temporal Logic Task Decomposition in Multi-Agent Systems

### v2

@@ 2.2 Signal Temporal Logic
where $G,F$ and $U$ are the temporal always, eventually and until operators with time interval $a,b]\subseteq\mathbb{R}_{+}$, while $\land,\lor$ and $\neg$ represent the logical conjunction, disjunction and negation operators. The qualitative semantic rules defining the conditions under which a state signal $\bm{x}(t)$ satisfies a task $\phi$ from time $t$ (written as $(\bm{x}(t),t)\models\phi$) were derived in [1, Ch 2.2]. Namely, the semantics relevant for the presentation are :
@@
$\varphi_{ij}:=F_{[a,b]}\mu_{ij}\;|\;G_{[a,b]}\mu_{ij},\\ ==$ (11b)
$\phi_{ij}:=\bigwedge^{K_{ij}}_{k=1}\varphi^{k}_{ij},\quad\phi_{i}:=\bigwedge^{K_{i}}_{k=1}\varphi^{k}_{i},$ (11c)
@@
for some centres $\bm{c}_{i},\bm{c}_{ij}\in\mathbb{R}^{n}$. It is assumed that the sets $\mathcal{B}_{i},\mathcal{B}_{ij}$ for $h_{i}(\bm{x}_{i})$, $h_{ij}(\bm{e}_{ij})$ are bounded, thus taking the form
@@
We refer to tasks of type (11a) and (11b) as independent tasks and collaborative tasks, respectively. Moreover, we refer to $\mathcal{B}_{ij}$ and $\mathcal{B}_{i}$ as the truth sets for tasks $\varphi_{ij}$ and $\varphi_{i}$ respectively since, by definition, these are the sets where $\mu_{ij}=\top$ and $\mu_{i}=\top$. In other words, independent and collaborative tasks enforce the single state and relative state of a pair of agents to lie within a bounded polytope for a time interval specified by the temporal operator included in the task. Note that recurrent tasks with period $T$ can be encoded in the fragment as conjunctions of the form $\land_{k}G_{[a+kT,b+kT]}\mu$. Safety specifications (e.g., collision avoidance) are not part (11), as we assume safety is enforced by the controller applied to satisfy the tasks.
@@
then considering, for convenience, the initial time $t=0$, we have the satisfaction relation $\rho^{\phi}(\bm{x}(t),0)\geq 0\;\Leftrightarrow\;(\bm{x}(t),0)\models\phi$ (with double implication holding due to absence of the negation operator in fragment (11)) [36, Sec. 2.2][37].
2.3 Communication and task graphs
Let $\mathcal{G}(\mathcal{V},\mathcal{E})$ denote an undirected graph over the agents in $\mathcal{V}$ with undirected edge set $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ such that $(i,j)\in\mathcal{E}\Leftrightarrow(j,i)\in\mathcal{E}$. Furthermore, let the neighbour set for agent $i$ be defined as $\mathcal{N}(i)=\{j\,|\,(i,j)\in\mathcal{E}\land i\neq j\}$. In the following, we consider two main graphs: the communication graph $\mathcal{G}_{c}(\mathcal{V},\mathcal{E}_{c})$ and the task graph $\mathcal{G}_{\psi}(\mathcal{V},\mathcal{E}_{\psi})$, such that
$\mathcal{E}_{c}$ $=\{(i,j)|\;\|\bm{p}_{i}-\bm{p}_{j}\|\leq r_{c}\}$ (16a)
@@ 2.3 Communication and task graphs
where $r_{c}>0$ is the communication radius. Note that by (16b) an edge $(i,j)\in\mathcal{E}_{\psi}$ exists whether a collaborative task $\phi_{ij}$ as per (11c) exists over the relative state $\bm{e}_{ij}$, while self-loops $(i,i)\in\mathcal{E}_{\psi}$ are induced by independent tasks $\phi_{i}$. Let $\mathcal{N}_{c}(i)$ and $\mathcal{N}_{\psi}(i)$ be the neighbour sets for the communication and task graphs respectively. When clear from the context, the shorthand notation $\mathcal{G}_{c}$ and $\mathcal{G}_{\psi}$ applies. In general, the communication graph is time-varying according to our definition since the position of the agents changes over time. However, we assume a supervisory controller enforcing the connectivity for each edge $(i,j)\in\mathcal{E}_{c}$ exists, such that the graph $\mathcal{G}_{c}$ is considered static (time-invariant) and determined at the initial time [38]. The following definition clarifies when a collaborative task is considered consistent with the communication requirements.
@@
On the other hand, $\varphi_{ij}$ is communication inconsistent if it is not consistent. Moreover, $\mathcal{G}_{\psi}$ is communication consistent if all collaborative tasks $\varphi_{ij}^{k},\;\forall(i,j)\in\mathcal{E}_{\psi},\;k=1,\ldots K_{ij}$, are communication consistent.
Relation (17) entails the existence of a nonempty set of relative configurations $\bm{e}_{ij}\in\mathbb{X}_{ij}$ whereby the predicate $\mu_{ij}$, associated with task $\varphi_{ij}$, takes the value $\mu_{ij}=\top$ while agents $i$ and $j$ are within the communication radius. This ensures the satisfiability of $\varphi_{ij}$ subject to the limited-range communication constraint. For brevity, we refer to communication consistent tasks simply as consistent.
2.4 Paths and cycles
For a given graph $\mathcal{G}(\mathcal{V},\mathcal{E})$, let the vector $\bm{\pi}_{i}^{j}\in\mathcal{V}^{l}$ represent a directed path of length $l$ defined as a vector of non-repeated indices in $\mathcal{V}$ such that $\bm{\pi}_{i}^{j}[k]\in\mathcal{V},\;\forall k=1,\ldots l$ ; $\bm{\pi}_{i}^{j}[k]\neq\bm{\pi}_{i}^{j}[k+1],\;\forall k=1,\ldots l-1$ ; $(\bm{\pi}_{i}^{j}[k],\bm{\pi}_{i}^{j}[k+1])\in\mathcal{E}$ and $(\bm{\pi}_{i}^{j}[1],\bm{\pi}_{i}^{j}[l])=(i,j)$. Similarly, let $\bm{\omega}\in\mathcal{V}^{l}$ represent a cycle such that $\bm{\omega}$ is a path of length $l$ with a single repeated index as $\bm{\omega}[1]=\bm{\omega}[l]$. A graph is acyclic when it does not contain cycles. Moreover, let $\epsilon:\mathcal{V}^{l}\rightarrow 2^{\mathcal{E}}$ be a set-valued function yielding the edges along the path $\bm{\pi}_{i}^{j}$ as $\epsilon(\bm{\pi}_{i}^{j})=\{(\bm{\pi}_{i}^{j}[k],\bm{\pi}_{i}^{j}[k+1])\;|\;k=1,\ldots l-1\}$, with the same definition extending to cycles $\bm{\omega}$. Thus, we can write
@@ 2.4 Paths and cycles
The following assumption over $\mathcal{G}_{c}$ is further considered.
Assumption 1
The communication graph $\mathcal{G}_{c}$ is connected and acyclic.
Assumption 1 is an initialization assumption imposing acyclicity of the communication graph $\mathcal{G}_{c}$. Note that when the communication graph is not naturally acyclic based on the communication range, then we assume that the agents can select a subset of the neighbours $\mathcal{N}_{c}(i)$ that renders $\mathcal{G}_{c}$ acyclic by electing a leader agent from which an acyclic graph can be constructed using standard graph search algorithms [39, Ch. 5]. Our previous work in [21], dealing with a similar decomposition, considered the milder assumption of connectivity for $\mathcal{G}_{c}$. The reason why Assumption 1 is considered here instead, relates to the fact that the convexity and decentralization properties of the task decomposition approach proposed in this work may be lost when cycles are present over $\mathcal{G}_{c}$. Further details justifying this assumption are given in Sec. 5.2.
3 Problem Formulation
@@
Global tasks as per (19) are suitable for defining time-varying relative formations from which complex group behaviours for the MAS can be obtained. The problem we approach is formally stated as follows:
@@
for all inconsistent edges $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ and with tasks $(\bar{\varphi}^{ij}_{rs})^{k},\;\forall(r,s)\in\epsilon(\bm{\pi}_{i}^{j}),\forall k=1,\ldots K_{ij}$ being communication consistent and as per (11b). Moreover, $\bar{\psi}$ should respect
@@
Section 5.1 details the design of tasks $\bar{\phi}^{ij}$. Before that, the next section presents key results to explain how these tasks can be introduced into the task graph $\mathcal{G}_{\psi}$ without inducing conflicts with other tasks in $\mathcal{G}_{\psi}$.
Figure 2: Decomposition of the inconsistent task $\phi_{13}=\varphi_{13}^{1}\land\varphi_{13}^{2}$ over the path $\bm{\pi}_{i}^{j}=[1,2,3]$. Solid lines show the communication graph $\mathcal{G}_{c}$; the dashed edge $(1,3)$ is the only one in the task graph $\mathcal{G}_{\psi}$.
4 Conflicting Conjunctions
@@
It is desired that any formula $\bar{\psi}$ as per (20), resulting from the decomposition approach, should not induce conflicting conjunctions as this fact would result directly in $\bar{\psi}$ being unsatisfiable. We here present five types of conflicting conjunctions according to Def. 5 which can arise when collaborative tasks are set in conjunction. Due to space limitations, proofs of the results in this section can be found in [34]. Intuitively, conflicting conjunctions of collaborative tasks in the fragment (11) occur when two or more collaborative tasks feature an intersection in the temporal domain (in terms of the intersection of their time intervals) which does not correspond to an intersection in the spatial domain (in terms of their truth sets). We here propose a more general definition of conflicting conjunction compared to the one in [21]. We conjecture that these are the only five types of conflicting conjunctions over collaborative tasks according to fragment (11). We acknowledge that conflicting conjunctions can also arise among conjunctions of independent and collaborative tasks, but we limit the presentation to conflicts over collaborative tasks. Adding conditions for conflicting conjunctions of independent and collaborative tasks is the subject of future work. In the presentation of the different types of conflicting conjunctions, we consider the general conjunction of tasks $\phi_{ij}=\bigwedge_{k=1}^{K_{ij}}\varphi^{k}_{ij}$ over an edge $(i,j)$ as per (11c), with truth sets $\mathcal{B}_{ij}^{k},\forall k=1,\ldots K_{ij}$ as per (14). Moreover, let the set of indices $\mathcal{I}_{G}$, $\mathcal{I}_{F}$ such that $\mathcal{I}_{G}\cup\mathcal{I}_{F}=\{1,\ldots K_{ij}\}$ with
@@
Intuitively, conflicting conjunctions as per Fact 1 occur any time a set of collaborative tasks, featuring an always operator, also feature a non-empty time interval intersection. Then, for any time in this time interval intersection, the relative state $\bm{e}_{ij}$ should lie inside the truth set $\mathcal{B}_{ij}$ of all the considered tasks, which is possible only if the truth sets are intersecting as well. The next two facts involve conflicting conjunctions of tasks with mixed temporal operators.
@@
Figure 3: Example of two conjunctions of tasks that do not induce conflicting tasks as per Fact 2. Detailed explanation in Example 2.
We provide some intuitions over the results of Facts 2 and 3. Consider a collaborative task $\varphi_{ij}^{d}=F_{[a^{d},b^{d}]}\mu_{ij}^{d}$, with $d\in\mathcal{I}_{F}$, such that its satisfaction can potentially occur at any time in the interval $a^{d},b^{d}]$ as per (LABEL:eq:eventually_old_def). Conflicts as per Fact 2 can then occur if there exists a set of tasks $\phi_{ij}^{l}=G_{[a^{l},b^{l}]}\mu_{ij}^{l}$ such that the union $\bigcup_{l}[a^{l},b^{l}]$ is covering the interval $a^{d},b^{d}]$ in the sense that $\bigcup_{l}[a^{l},b^{l}]\supseteq[a^{d},b^{d}]$. In this case at least one of the tasks $\phi_{ij}^{l}$ must necessarily be satisfied together with $\varphi_{ij}^{d}$. On the other hand, conflicts as per Fact 3 occur when the time interval intersection of the tasks $\varphi_{ij}^{l}$ is covering $a^{d},b^{d}]$ in the sense that $\bigcap_{l}[a^{l},b^{l}]\supseteq[a^{d},b^{d}]$. In this case, there exists at least one time instant $\tau\in[a^{d},b^{d}]$ such that all tasks $\phi_{ij}^{l}=G_{[a^{l},b^{l}]}\mu_{ij}^{l}$ must necessarily be satisfied together with $\varphi_{ij}^{d}=F_{[a^{d},b^{d}]}\mu_{ij}^{d}$. An example is provided to support these proposed intuitions.
@@
While we provided the statement for conflicting conjunction of type 4 and 5, we show in Section 5.2 that these can not arise for the proposed decomposition approach due to the acyclicity assumption on $\mathcal{G}_{c}$. We choose to report them as they represent an independent contribution to future work in this direction. The following assumption is considered further
@@
This is a reasonable assumption over $\mathcal{G}_{\psi}$ as otherwise the task $\psi$ would be unsatisfiable even before the decomposition. Section 5.2 details how the proposed decomposition approach preserves the conflict-free property passing from $\mathcal{G}_{\psi}$ to $\mathcal{G}_{\bar{\psi}}$.
5 Task Decomposition
In this section, the decomposition approach to obtain tasks $\bar{\phi}^{ij}$ from the inconsistent tasks $\phi_{ij}$ defined over the edges $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ is presented as per Problem 1. For the sake of clarity and to reduce the burden of notation, we present our results assuming that all the inconsistent tasks $\phi_{ij}$ do not contain conjunctions such that simply $\phi_{ij}=\varphi_{ij}$. Hence, by dropping the index $k$ in (21) and replacing (21b) in (21a), we get that $\phi_{ij}=\varphi_{ij}$ is decomposed into the task $\bar{\phi}^{ij}=\land_{(r,s)\in\epsilon(\bm{\pi}_{i}^{j})}\bar{\varphi}_{rs}^{ij}$. Note that this simplifying assumption is not required from our approach but is nonetheless introduced for the sake of clarity. Indeed, the same approach developed to decompose a single collaborative task $\varphi_{ij}$ into an appropriate task $\bar{\phi}^{ij}$ as per (21b), can be repeated for any index $k\in\{1,\ldots K_{ij}\}$ in the conjunction. With these considerations, in Sec. 5.1 the first main decomposition result is provided in the form of Lemma 1, where a procedure to define the task $\bar{\phi}^{ij}$ from an inconsistent task $\varphi_{ij}$ is derived. The task $\bar{\phi}^{ij}$ resulting from this decomposition is a function of a set of parameters that are later shown to be computable via convex optimization. Moreover, Sec. 5.2 and 5.3 detail how convex constraints over the former parameters are enforced to avoid conflicting conjunctions and guarantee communication consistency.
5.1 Parametric tasks for decomposition
Consider the communication inconsistent task $\phi_{ij}=\varphi_{ij}$ with $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$, where $h_{ij}=\langle-(A_{ij}(\bm{e}_{ij}-\bm{c}_{ij})-\bm{z}_{ij})\rangle$ is the predicate function associated with $\varphi_{ij}$ and $\mathcal{B}_{ij}=\mathcal{P}(\bm{A}_{ij},\bm{c}_{ij},\bm{z}_{ij})$ is the associated truth set as in (13)-(14). Furthermore, consider the family of parametric predicate functions

### v3

@@ II-B Signal Temporal Logic
where $G,F$ and $U$ are the temporal always, eventually and until operators with time interval $a,b]\subseteq\mathbb{R}_{+}$, while $\land,\lor$ and $\neg$ represent the logical conjunction, disjunction and negation operators. The qualitative semantic rules defining the conditions under which a state signal $\bm{x}(t)$ satisfies a task $\phi$ from time $t$ (written as $(\bm{x}(t),t)\models\phi$) are as follows [1, Ch 2.2]:
[…]
@@
$\varphi_{ij}:=F_{[a,b]}\mu_{ij}|G_{[a,b]}\mu_{ij},\\$ (12b)
$\phi_{ij}:=\bigwedge^{K_{ij}}_{k=1}\varphi^{k}_{ij},\;\phi_{i}:=\bigwedge^{K_{i}}_{k=1}\varphi^{k}_{i},$ (12c)
@@
for some centres $\bm{c}_{i},\bm{c}_{ij}\in\mathbb{R}^{n}$. It is assumed that the super-level sets $\mathcal{B}_{i},\mathcal{B}_{ij}$ for $h_{i}(\bm{x}_{i})$, $h_{ij}(\bm{e}_{ij})$ are bounded, thus taking the form
@@
We refer to tasks of type (12a) and (12b) as independent tasks and collaborative tasks respectively. Moreover, we refer to $\mathcal{B}_{ij}$ and $\mathcal{B}_{i}$ as the truth sets for tasks $\varphi_{ij}$ and $\varphi_{i}$ respectively since, by definition, these are the sets where $\mu_{ij}=\top$ and $\mu_{i}=\top$. In other words, independent and collaborative tasks enforce the single state and relative state of a pair of agents to lie within a bounded polytope for a time interval specified by the temporal operator included in the task. Note that (12) is expressive enough to deal with tasks of type $\mu^{1}U_{[a,b]}\mu^{2}$ since $\mu^{1}U_{[a,b]}\mu^{2}=G_{[a,\tau]}\mu^{1}\land F_{[\tau,\tau]}\mu^{2}$ with $\tau\in[a,b]$. At the same time, recurrent tasks with period $T$ can be encoded in the fragment as conjunctions of the form $\land_{k}G_{[a+kT,b+kT]}\mu$.
@@
then considering for convenience the initial time $t=0$ we have the satisfaction relation $\rho^{\phi}(\bm{x}(t),0)\geq 0\;\Leftrightarrow\;(\bm{x}(t),0)\models\phi$ (with double implication holding due to absence of the negation operator in fragment (12)) [35, Sec. 2.2][36].
Remark 1
While generally nonlinear concave predicate functions are commonly considered in the literature, predicate functions with polytope-like truth sets as in (14) represent a practical design choice. Indeed, let any two predicate functions $h_{1}(\bm{x}):\mathcal{D}\rightarrow\mathbb{R}$ and $h_{2}(\bm{x}):\mathcal{D}\rightarrow\mathbb{R}$ such that the truth set $\mathcal{B}=\{\bm{x}\;|\;h_{1}(\bm{x})\geq 0\}=\{\bm{x}\;|\;h_{2}(\bm{x})\geq 0\}$ then $\mu(h_{1}(\bm{x}))=\mu(h_{2}(\bm{x})),\,\forall\bm{x}\in\mathcal{D}$. Hence, different predicate functions with the same truth set are semantically indistinguishable as per the semantics in (11) and (16). Since any smooth convex set $\mathcal{B}$ can be approximated by a polytope $\mathcal{P}(A,\bm{c},\bm{z})$, with the approximation error decreasing in the number of hyperplanes used for the approximation, predicate functions as per (14) are considered comparable to nonlinear concave predicate functions, with the critical advantage that set operations like Minkowski sums, set inclusions and intersections can be efficiently undertaken [37][38, Ch. 1.10].
II-C Communication and task graphs
Let $\mathcal{G}(\mathcal{V},\mathcal{E})$ represent an undirected graph over the agents in $\mathcal{V}$ with undirected edge set $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ such that $(i,j)\in\mathcal{E}\Leftrightarrow(j,i)\in\mathcal{E}$. Furthermore, let the neighbour set for agent $i$ be defined as $\mathcal{N}(i)=\{j\,|\,(i,j)\in\mathcal{E}\land i\neq j\}$. In the following, a distinction is made between two types of graphs: the communication graph $\mathcal{G}_{c}(\mathcal{V},\mathcal{E}_{c})$ and the task graph $\mathcal{G}_{\psi}(\mathcal{V},\mathcal{E}_{\psi})$, such that
$\mathcal{E}_{c}$ $=\{(i,j)|\;\|\bm{p}_{i}-\bm{p}_{j}\|\leq r_{c}\land q^{c}_{ij}=\top\}$ (17a)
@@ II-C Communication and task graphs
In (17a), $r_{c}>0$ is the communication radius and $q^{c}_{ij}\in\{\top,\bot\},\;\forall(i,j)\in\mathcal{V}\times\mathcal{V}$ is a boolean communication token that is applied to selectively enable communication among agents that are within the communication radius. For consistentency with self-communication, let $q^{c}_{ii}=\top,\;\forall i\in\mathcal{V}$. On the other hand, note that by (17b) an edge $(i,j)\in\mathcal{E}_{\psi}$ exists whether a collaborative task $\phi_{ij}$ as per (12c) exists over the relative state $\bm{e}_{ij}$, while self-loops $(i,i)\in\mathcal{E}_{\psi}$ are induced by independent tasks $\phi_{i}$. Let $\mathcal{N}_{c}(i)$ and $\mathcal{N}_{\psi}(i)$ be the neighbour sets for the communication and task graphs respectively. When clear from the context, the shorthand notation $\mathcal{G}_{c}$ and $\mathcal{G}_{\psi}$ applies. In general, the communication graph is time-varying according to our definition since the position of the agents changes over time. However, we assume a supervisory controller enforcing the connectivity for each edge $(i,j)\in\mathcal{E}_{c}$ exists, such that the graph $\mathcal{G}_{c}$ is considered static (time-invariant) and determined at the initial time [39]. The following definition clarifies when a collaborative task is considered consistent with the communication requirements.
@@
On the other hand, $\varphi_{ij}$ is communication inconsistent if it is not consistent. Moreover, $\mathcal{G}_{\psi}$ is communication consistent if all the collaborative tasks $\varphi_{ij}^{k},\;\forall(i,j)\in\mathcal{E}_{\psi}\;k\in$, are.
Communication consistency as per (4) plays a central role in our task decomposition approach. Namely, relation (18) entails that there exists a nonempty set of relative configuration $\bm{e}_{ij}\in\mathbb{X}_{ij}$ such that the predicate $\mu_{ij}$, associated with task $\varphi_{ij}$, takes the value $\mu_{ij}=\top$ while agent $i$ and $j$ are within the communication radius. This fact thus ensures the satisfiability of $\varphi_{ij}$ subject to the limited-range communication constraint. For brevity, we will often refer to communication consistent tasks simply as consistent without loss of clarity. The following assumption about the communication graphs is considered.
Assumption 1
The communication graph $\mathcal{G}_{c}$ is connected and the communication tokens $q_{ij}^{c}$ are chosen such that $\mathcal{G}_{c}$ is acyclic.
Assumption 1 is an initialization assumption imposing acyclicity of the communication graph $\mathcal{G}_{c}$. Our previous work in [21], dealing with a similar decomposition, considered the milder assumption of connectivity for $\mathcal{G}_{c}$. The reason why Assumption 1 is considered here instead, relates to the fact that the convexity and decentralization property of the task decomposition approach proposed in the work are lost when cycles are present over $\mathcal{G}_{c}$. Further details justifying this assumption are given in Sec. V-B. We highlight that the acyclicity of $\mathcal{G}_{c}$ can be enforced by appropriately selecting the communication tokens $q_{ij}^{c}$ at time $t=0$ by means of centralized or decentralized algorithms. One such example consists of electing a root agent and expanding a tree that spans all the agents similar to well-known sampling-based planning algorithms over graphs [40, Ch. 5].
II-D Paths, cycles and critical tasks
For a given graph $\mathcal{G}(\mathcal{V},\mathcal{E})$, let the vector $\bm{\pi}_{i}^{j}\in\mathcal{V}^{l}$ represent a directed path of length $l$ defined as a vector of non-repeated indices in $\mathcal{V}$ such that $\bm{\pi}_{i}^{j}[k]\in\mathcal{V},\;\forall k=1,\ldots l$ ; $\bm{\pi}_{i}^{j}[k]\neq\bm{\pi}_{i}^{j}[k+1],\;\forall k=1,\ldots l-1$ ; $(\bm{\pi}_{i}^{j}[k],\bm{\pi}_{i}^{j}[k+1])\in\mathcal{E}$ and $(\bm{\pi}_{i}^{j}[1],\bm{\pi}_{i}^{j}[l])=(i,j)$. Similarly, let $\bm{\omega}\in\mathcal{V}^{l}$ represent a cycle such that $\bm{\omega}$ is a path of length $l$ with a single repeated index as $\bm{\omega}[1]=\bm{\omega}[l]$. Moreover, let $\epsilon:\mathcal{V}^{l}\rightarrow 2^{\mathcal{E}}$ be a set-valued function yielding the edges along the path $\bm{\pi}_{i}^{j}$ as $\epsilon(\bm{\pi}_{i}^{j})=\{(\bm{\pi}_{i}^{j}[k],\bm{\pi}_{i}^{j}[k+1])\;|\;k=1,\ldots l-1\}$, with the same definition extending to cycles $\bm{\omega}$. From these definitions, the following relations hold
@@ II-D Paths, cycles and critical tasks
III Problem Formulation
@@
Global tasks as per (20) are suitable for defining time-varying relative formations from which complex group behaviours for the MAS can be obtained. Previously developed feedback-based control laws to satisfy STL tasks in these settings take the general form $\bm{u}_{i}(t)=\bm{u}_{i}(\bm{x}_{i}(t),\bm{x}_{\mathcal{N}_{c}(i)}(t),t)$ with $\bm{x}_{\mathcal{N}_{c}(i)}(t)=[\bm{x}_{i}(t)]_{i\in\mathcal{N}_{c}(i)}$ [18, 41, 26, 27]. However, a complete [^2] A graph is complete if every agent is connected to every other agent. communication graph $\mathcal{G}_{c}$ is implicitly assumed such that $\bm{x}_{\mathcal{N}_{c}(i)}(t)=[\bm{x}_{i}(t)]_{i\in\mathcal{V}}$. More recently in [42], we provided a feedback control approach for the satisfaction of global tasks in the form (20), assuming that the associated task graph and communication graph are consistent as well as acyclicity of the task graph, which we also assume in the current work. This fact motivates the effort for a decomposition approach for rewriting communication inconsistent tasks $\varphi_{ij}$ (cf. Def. 4) as conjunctions of consistent ones, such that the advantages of state feedback can be leveraged to achieve the satisfaction of the task. The problem is formally stated as follows:
@@
for all inconsistent edges $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ and with tasks $(\bar{\varphi}^{ij}_{rs})^{k},\;\forall(r,s)\in\epsilon(\bm{\pi}_{i}^{j}),\forall k=1,\ldots K_{ij}$ being communication consistent as per (12b). Moreover, $\bar{\psi}$ should be such that
@@
In Section V-A it is detailed how the tasks $\bar{\phi}^{ij}$ are designed. But before, in the next section, some important results are presented to understand the process under which the tasks $\bar{\phi}^{ij}$ can be introduced over the task graph $\mathcal{G}_{\psi}$ without causing conflicts.
Figure 2: Graphical decomposition of the inconsistent task $\phi_{13}=\varphi_{13}^{1}\land\varphi_{13}^{2}$ over the path $\bm{\pi}_{i}^{j}=[1,2,3]$. Solid edges correspond to edges in the communication graph $\mathcal{G}_{c}$ and edge $(1,3)$ (dashed) is the only edge in the task graph $\mathcal{G}_{\psi}$.
IV Conflicting conjunctions over task graphs
@@
It is desired that any formula $\bar{\psi}$ as per (21), resulting from the decomposition approach, should not induce conflicting conjunctions as this fact would result directly in $\bar{\psi}$ being unsatisfiable. We here study five types of conflicting conjunctions according to Def. 5 which can arise when collaborative tasks are set in conjunction. Proofs of the results are given in Appendix A.
Intuitively, conflicting conjunctions of collaborative tasks in the fragment (12) occur when two or more collaborative tasks feature an intersection in the temporal domain (in terms of the intersection of their time intervals) which does not correspond to an intersection in the spatial domain (in terms of their truth sets). We here propose a more general definition of conflicting conjunction compared to the one in [21]. We conjecture that these are the only five types of conflicting conjunctions over collaborative tasks according to fragment (12). We acknowledge that conflicting conjunctions can also arise among conjunctions of independent and collaborative tasks, but we limit the presentation to conflicts over collaborative tasks. Adding conditions for conflicting conjunctions of independent and collaborative tasks is the subject of future work.
In the presentation of the different types of conflicting conjunctions, we consider the general conjunction of tasks $\phi_{ij}=\bigwedge_{k=1}^{K_{ij}}\varphi^{k}_{ij}$ over an edge $(i,j)$ as per (12c), with truth sets $\mathcal{B}_{ij}^{k},\forall k=1,\ldots K_{ij}$ as per (15). Moreover, let the set of indices $\mathcal{I}_{G}$, $\mathcal{I}_{F}$ such that $\mathcal{I}_{G}\cup\mathcal{I}_{F}=\{1,\ldots K_{ij}\}$ with
@@
Concerning the intuitive meaning of conflicting conjunctions as per Fact 1, these occur any time a set of collaborative tasks, featuring an always operator, also feature a non-empty time interval intersection. Indeed, for any time in this time interval intersection, the relative state $\bm{e}_{ij}$ should be inside the truth set $\mathcal{B}_{ij}$ of all the considered tasks, which is possible only if the truth sets are intersecting as well. The next two facts involve conflicting conjunctions of tasks with mixed temporal operators.
@@
Figure 3: Graphical exemplification of two conjunctions of tasks that do not induce conflicting tasks as per Fact 2. Detailed explanation in Example 2.
We provide some intuitions over the results of Facts 2 and 3. Consider a collaborative task $\varphi_{ij}^{d}=F_{[a^{d},b^{d}]}\mu_{ij}^{d}$, with $d\in\mathcal{I}_{F}$, such that its satisfaction can potentially occur at any time in the interval $a^{d},b^{d}]$ as per (LABEL:eq:eventually_old_def). Conflicts as per Fact 2 can then occur if there exists a set of tasks $\phi_{ij}^{l}=G_{[a^{l},b^{l}]}\mu_{ij}^{l}$ such that the union $\bigcup_{l}[a^{l},b^{l}]$ is covering the interval $a^{d},b^{d}]$ in the sense that $\bigcup_{l}[a^{l},b^{l}]\supseteq[a^{d},b^{d}]$. In this case at least one of the tasks $\phi_{ij}^{l}$ must necessarily be satisfied together with $\varphi_{ij}^{d}$. On the other hand, conflicts as per Fact 3 occur when the time interval intersection of the tasks $\varphi_{ij}^{l}$ is covering $a^{d},b^{d}]$ in the sense that $\bigcap_{l}[a^{l},b^{l}]\supseteq[a^{d},b^{d}]$. In this case, there exists at least one time instant $\tau\in[a^{d},b^{d}]$ such that all tasks $\phi_{ij}^{l}=G_{[a^{l},b^{l}]}\mu_{ij}^{l}$ must necessarily be satisfied together with $\varphi_{ij}^{d}=F_{[a^{d},b^{d}]}\mu_{ij}^{d}$. The next example provides a graphical explanation for these intuitions.
@@
While we provided the statement for conflicting conjunction of type 4 and 5, we show in Section V-B that these can not arise during our proposed decomposition thanks to the acyclicity assumption over the communication graph $\mathcal{G}_{c}$. We choose to report them as they represent an independent contribution to future work in this direction. The following assumption is considered further
@@
This is a reasonable assumption over $\mathcal{G}_{\psi}$ since otherwise the task $\psi$ would be unsatisfiable even before the decomposition. We explore in Section V-B how the proposed decomposition approach preserves the conflict-free property passing from $\mathcal{G}_{\psi}$ to $\mathcal{G}_{\bar{\psi}}$.
V Task Decomposition
In this section, the decomposition approach to obtain tasks $\bar{\phi}^{ij}$ from the inconsistent tasks $\phi_{ij}$ defined over the edges $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ is presented as per Problem 1. For the sake of clarity and to reduce the burden of notation, we present our results assuming that all the inconsistent tasks $\phi_{ij}$ do not contain conjunctions such that simply $\phi_{ij}=\varphi_{ij}$. Hence, by dropping the index $k$ in (22) and replacing (22b) in (22a), we get that $\phi_{ij}=\varphi_{ij}$ is decomposed into the task $\bar{\phi}^{ij}=\land_{(r,s)\in\epsilon(\bm{\pi}_{i}^{j})}\bar{\varphi}_{rs}^{ij}$. Note that this simplifying assumption is not required from our approach but is nonetheless introduced for the sake of clarity. Indeed, the same approach developed to decompose a single collaborative task $\varphi_{ij}$ into an appropriate task $\bar{\phi}^{ij}$ as per (22b), can be repeated for any index $k\in\{1,\ldots K_{ij}\}$ in the conjunction. With these considerations, in Sec. V-A the first main decomposition result is provided in the form of Lemma 1, where a procedure to define the task $\bar{\phi}^{ij}$ from an inconsistent task $\varphi_{ij}$ is derived. The task $\bar{\phi}^{ij}$ resulting from this decomposition is a function of a set of parameters that are later shown to be computable via convex optimization. Moreover, Sec. V-B and V-C detail how convex constraints over the former parameters can be enforced such that conflicting conjunctions are avoided and each task is communication consistent.
V-A Parametric tasks for decomposition
Consider the communication inconsistent task $\phi_{ij}=\varphi_{ij}$ with $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$, where $h_{ij}=\langle-(A_{ij}(\bm{e}_{ij}-\bm{c}_{ij})-\bm{z}_{ij})\rangle$ is the predicate function associated with $\varphi_{ij}$ and $\mathcal{B}_{ij}=\mathcal{P}(\bm{A}_{ij},\bm{c}_{ij},\bm{z}_{ij})$ is the associated truth set as per (14) and (15). Furthermore, consider the following family of parametric predicate functions